### v1

@@ 5.2 Performance of the dynamic models
Table 5 and Table 6 show that, for each of the single layer networks, the GAT-LSTM-ATT model produces the highest AUC and $F_{1}$ scores, while GCN-GRU gives the poorest results. This could be due to the fact that GAT assigns different importance to each edge, and we know that some connections could be more informative than others. Also, the complex structure of LSTM appears to make it the preferred RNN for this problem. Another key observation is that models enhanced with the attention mechanism consistently show better performance compared to those without attention.
@@
As for the double layer network, we can see from Table 7 that, similarly to what was observed for the single layer networks, the GAT-LSTM-ATT model again shows the best performance. The results for the double layer network, however, tend to outperform the single layer ones, which is intuitive as the double layer network is able to consider connections of either type. It is also noticeable that the results obtained from the double layer network have shorter confidence intervals, suggesting greater robustness in these results. Importantly, the best performing dynamic modelling approach, i.e., GAT-LSTM-ATT, consistently outperforms any of the baseline methods presented in the previous subsection, often by a considerable margin. For example, the GAT-LSTM-ATT model for the double-layer network produces AUC and $F_{1}$ scores of 0.812 and 0.851, respectively, compared to 0.805 and 0.837 for the best baseline model, i.e., XGB (see Table 4). Note that this holds not just for the double layer network, but for either of the single layer networks as well. Hence, incorporating dynamic network information is able to provide additional information over simply using local features and/or static networks. The most pronounced difference lies between our dynamic model and the static network-based models, demonstrating the importance of capturing network changes over a sufficiently long time window.
@@ 5.3 Runtime analysis
From the tables, we can see that training a model on a single layer network derived from the lending company takes longer than training a model on a single layer network created based on geographical area. This is not unexpected as the former network contains a higher number of connections between the nodes compared to the latter. GAT-LSTM and GAT-LSTM-ATT have the highest training runtimes among the GNN-RNN and GNN-RNN-ATT models, respectively, whereas GCN-GRU and GCN-GRU-ATT have the lowest runtimes. This is intuitive as GNN-RNN-ATT models normally have higher runtimes compared to the GNN-RNN models, owing to the complexity added to those models by the attention mechanism.
@@ 5.4 Interpretability of the architecture
(b) Average Shapley value for GAT-LSTM-ATT.
@@
The figure shows that for the first few timestamps, the attention score is relatively stable and low, remaining close to 0.1. This indicates a minimal level of attention or importance being assigned during these early timestamps. However, as time progresses, particularly after timestamp 3, there is a noticeable upward trend in the attention score. This increase becomes more pronounced between timestamps 4 and 6, where the attention score rises sharply, peaking just below 0.6 at timestamp 6. This pattern suggests that as time progresses, the snapshots grow in importance. The reason for this progressive increase in attention could be that the most recent information holds greater value.
@@ 6 Conclusions
The evaluation of these models was conducted using a dataset from the mortgage lending domain. In our experiments, we constructed single and double layer networks using the borrower’s geographical location and the lending company as the connector variables. Through rigorous testing of various models, we established that the GAT-LSTM-ATT model exhibits superior performance across all configurations of DYMGNN, also outperforming other baseline models, both GNN-based and non-GNN-based.
@@
Future research could explore wider networks by incorporating additional layers to map more complex inter-individual connections. It might also be beneficial to vary the number of snapshots by adjusting the window length for generating dynamic networks. The exploration of other GNNs and RNNs not considered in this study presents another promising direction. Moreover, gaining a deeper understanding of how different network connections influence default risk could offer valuable insights into credit risk modelling.

### v2

@@ 5.2 Performance of the dynamic models
Table 5 and Table 6 show that, for each of the single layer networks, the GAT-LSTM-ATT model produces the highest AUC and $F_{1}$ score, while GCN-GRU gives the poorest results. This could be due to the fact that GAT assigns different importance to each edge, and we know that some connections could be more informative than others. Also, the complex structure of LSTM appears to make it the preferred RNN for this problem. Another key observation is that models enhanced with the attention mechanism consistently show better performance compared to those without attention.
@@
As for the double layer network, we can see from Table 7 that, similarly to what was observed for the single layer networks, the GAT-LSTM-ATT model again shows the best performance. The results for the double layer network, however, tend to outperform the single layer ones, which is intuitive as the double layer network is able to consider connections of either type. It is also noticeable that the results obtained from the double layer network have shorter confidence intervals, suggesting greater robustness in these results. Importantly, the best performing dynamic modelling approach, i.e., GAT-LSTM-ATT, performs better on average than the baseline methods presented in the previous subsection. For example, the GAT-LSTM-ATT model for the double-layer network produces AUC and $F_{1}$ score of 0.812 and 0.851, respectively, compared to 0.805 and 0.837 for the best baseline model, i.e., XGB (see Table 4). This translates to a 0.87% gain in AUC and a 1.67% gain in $F_{1}$ score. Although these numerical gains might seem modest, even a 1% improvement can yield significant financial benefits for some businesses. Interestingly, even when applied to either of the single layer networks, the GAT-LSTM-ATT still tends to perform well compared to the baseline models. This demonstrates that DYMGNN offers an advantage over conventional methods by capturing a richer set of information, thus providing a more comprehensive and realistic picture of a borrower’s default probability. Hence, incorporating dynamic network information is able to provide additional information over simply using local features and/or static networks. The most pronounced difference lies between our dynamic model and the static network-based models, demonstrating the importance of capturing network changes over a sufficiently long time window.
@@ 5.3 Runtime analysis
From the tables, we can see that training a model on a single layer network derived from the lending company takes longer than training a model on a single layer network created based on geographical area. This is not unexpected as the former network contains a higher number of connections between the nodes compared to the latter. GAT-LSTM and GAT-LSTM-ATT have the highest training runtimes among the GNN-RNN and GNN-RNN-ATT models, respectively, whereas GCN-GRU and GCN-GRU-ATT have the lowest runtimes. GNN-RNN-ATT models normally have higher runtimes compared to the GNN-RNN models, owing to the complexity added to those models by the attention mechanism. It is worth noting that the runtime for the XGB model is only 151 seconds, highlighting the greater complexity of our proposed models compared to traditional non-network classifiers.
@@ 5.4 Interpretability of the architecture
(b) Shapley values for GAT-LSTM-ATT.
@@
The figure shows that for the first few timestamps, the attention score is relatively stable and low, remaining close to 0.1. This indicates a minimal level of attention or importance being assigned during these early timestamps. However, as time progresses, particularly after timestamp 3, there is a noticeable upward trend in the attention score. This increase becomes more pronounced between timestamps 4 and 6, where the attention score rises sharply, peaking just below 0.6 at timestamp 6. This pattern suggests that as time progresses, the snapshots grow in importance. The reason for this progressive increase in attention could be that the most recent information holds greater value. Additionally, it can be inferred that longer lookback periods are unlikely to add anything extra, since the bulk of attention is allocated within a relatively brief period.
@@ 6 Conclusions
The evaluation of these models was conducted using a dataset from the mortgage lending domain. In our experiments, we constructed single and double layer networks using the borrower’s geographical location and the lending company as the connector variables. Through rigorous testing of various models, we established that the GAT-LSTM-ATT model exhibits the best performance among all configurations of DYMGNN, and other baseline models, both GNN-based and non-GNN-based. Furthermore, this model is able to capture a richer set of information, thereby providing more realistic insights into a borrower’s probability of default.
@@
Future research could explore wider networks by incorporating additional layers to map more complex inter-individual connections. The method could also be extended by further considering distance information and assigning different weights to the network edges based on geographical proximity between the centroids of neighbouring zip code areas. It might also be beneficial to vary the number of snapshots by adjusting the window length for generating dynamic networks. The exploration of other GNNs and RNNs not considered in this study presents another promising direction. Moreover, gaining a deeper understanding of how different network connections influence default risk could offer valuable insights into credit risk modelling.